# Robust LLM Alignment via Distributionally Robust Direct Preference Optimization

### v2

@@ Appendix C Proof of KLDPO Sample Complexity
Note that the function $g(\theta)=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex with respect to $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$ by Lemma 10. We use this fact in $(a)$. The inequality in $(b)$ is due to $\sup_{x}(f(x)+g(x))\leq\sup_{x}f(x)+\sup_{x}g(x)$. The last inequality $(c)$ is because $\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}})\geq\lambda$, for all $\mathsf{P}\in\mathcal{P}_{\mathrm{KL}}$ by Assumption 2. This implies that $\mathcal{L}^{\mathrm{KL}}$ is a $\gamma\lambda$-strongly convex function with respect to $\lVert\cdot\rVert_{2}$. ∎
@@ C.1 Proof of Policy Parameter Convergence of KLDPO
By Lemma 13, we have that, with probability at least $1-\delta$,
@@
where the first inequality is because $\theta^{\mathrm{KL}}_{n}$ is the minimizer of $\mathcal{L}^{\mathrm{KL}}_{n}$. Now by the $\gamma\lambda$-strong convexity of $\mathcal{L}^{\mathrm{KL}}$ (see Lemma 14) and Lemma 5.II, we have that
@@ Appendix D Proof of Tractable KLDPO
Next, we prove the formal version of Proposition 2.
@@
Here, $\mu$ and $\lambda$ are implicitly defined by the constraints (Eq. 24-Eq. 26). Now we prove that the dual variables $-\mu-\lambda$ can be upper bounded by $-\sum_{i=1}^{n}q_{i}l(z_{i};\theta)$.
@@
$-\mu-\lambda$ satisfies $-\mu-\lambda\leq-\sum_{i=1}^{n}\mathsf{P}_{n}^{o}(i)l(z_{i};\theta)$.
@@
This implies that $-\mu-\lambda\leq-\sum_{i=1}^{n}q_{i}l(z_{i};\theta)$. ∎
@@ E.1 ArmoRM Multi-objective Alignment
Similar to the Emotion Alignment experiments, we generate preference labels according to convex combinations of two reward objectives, i.e., $r^{*}_{\mathrm{convex}}$ defined as Mixture Evaluation in previous section. Specifically, we consider three pairs of objectives: (1) Ultrafeedback-Honesty and Helpsteer-Complexity, (2) Ultrafeedback-Helpfulness and Helpsteer-Coherence, and (3) Ultrafeedback-truthfulness and Helpsteer-Complexity. We generate preference labels according to $\alpha^{o}=0.5$ for all three cases. All models are trained for 4 epochs. Then we introduce reward shift by evaluating WDPO, KLDPO, and DPO on $r^{*}_{\mathrm{convex}}(\alpha)$, where $\alpha\in\{0,0.1,0.3,0.5,0.7,0.9,1\}$. In the first plot of Fig. 4, the training preferences are generated accoding to reward pair (1). We observe that WDPO and KLDPO achieve superior performance compared to DPO. In particular, when reward distribution shift happens in two directions (towards standalone Ultrafeedback-Honesty and Helpsteer-Complexity), they clearly outperform DPO. In the middle plot, the training preferences are generated according to reward pair (2). We observe that WDPO with both $\rho_{o}=0.005$ and $\rho_{o}=0.01$ are particularly robust against reward distribution shift. Lastly, in the third plot, the training preferences are generated according reward pair (3). We observe that both WDPO and KLDPO achieve notable robustness compared to DPO.
@@ E.2 Leaderboard Alignment
In this section, we include alignment results evaluated on all Leaderboard (Fourrier et al.,, 2024) sub-tasks. We list all sub-task names in Table 2.
Table 2: All sub-task names in OpenLLM Leaderboard 2.
@@
LLaMA-3.2-1B results: In Table 3, we compare DPO, KLDPO, and WDPO trained using LLaMA-3.2-1B on all 39 sub-tasks of OpenLLM Leaderboard. We observe that our WDPO and KLDPO methods achieve superior alignment performance on the majority of subtasks. Although WDPO and KLDPO slightly underperform on a few subtasks, their primary strength lies in generalization—precisely because they consistently enhance performance across a diverse range of subtasks.
LLaMA-3.1-8B results: In Table 4, we compare DPO and KLDPO, both trained using LLaMA-3.1-8B. Earlier, we demonstrated that WDPO, trained on LLaMA-3.2-1B, outperforms both DPO and KLDPO. However, WDPO’s requirement for dual gradient computations increases its computational complexity. Due to resource constraints, we present KLDPO results for the 8B model, as it is more scalable. Following the LLaMA-3.2-1B experiments, we train KLDPO for two epochs—the point where DPO achieved optimal robustness via early stopping. Notably, KLDPO exhibits exceptional performance on math-related tasks.
@@
Table 3: Evaluation of DPO, KLDPO, and WDPO on all OpenLLM Leaderboard 2 sub-tasks.
@@
Table 4: Evaluation of DPO and KLDPO on all OpenLLM Leaderboard 2 sub-tasks.
@@ Appendix F Additional Experiment Details
Reward Model Training: The raw Emotion dataset (Saravia et al.,, 2018) consists of text samples paired with multi-class labels for six different emotion classes (joy, sadness, love, anger, fear, and surprise). This dataset was then transformed into a multi-label dataset, referred to as the Emotion Reward Dataset. To create the multi-label dataset, the surprise class was excluded due to its limited representation in the original dataset. Following this, up to three random text samples from the raw dataset were concatenated, and their associated labels were merged. This pre-processing step ensured that the reward model encountered text samples representing multiple emotions during training.
@@
Supervised Fine-Tuning: Before training the WDPO algorithm, it is essential to ensure that the model familiarize with the types of texts present in the dataset. To achieve this, we performed supervised fine-tuning (SFT). We selected GPT-2 as the base language model and trained it to predict the next token based on the text samples in the emotion dataset. The maximum length of each text sample was capped at 68 tokens. The model was trained for 10 epochs with a batch size of 64. The training used the Adam optimizer (Kingma and Ba,, 2014) with a learning rate of $5.0\times 10^{-7}$ following 12 warmup steps. Additionally, a maximum gradient norm of 10 was applied to stabilize the training. The ArmoRM setups did not need any SFT as we used Intruct models which have already undergone multiple rounds of SFT and alignment.
@@
WDPO Implementation: (1) In WDPO training, one of the main challenges is calculating the gradient penalty of the DPO loss with respect to the input. However, since the input is tokenized as integers, gradient cannot be directly calculated. To address this, gradient is calculated with respect to the output of the embedding layer, where gradients are tracked. (2) In line 4 of the tractable WDPO algorithm (Algorithm 1), we compute the gradient regularizer: $\mathcal{R}(\pi_{\theta};\mathcal{D})=\rho_{o}(\mathbb{E}_{z\sim\mathcal{D}}\lVert\nabla_{z}l(z;\theta)\rVert_{2}^{2})^{1/2}$. A key implementation challenge arises in distributed LLM training. A naive approach computes the gradient of the pointwise DPO loss with respect to each input, averages the gradient norms over the micro-batch, and applies this as a regularizer to the batch DPO loss on each worker. However, due to the typically small micro-batch sizes in large-scale LLM training, this averaging is performed over very few samples, resulting in a highly noisy and unstable gradient penalty. To mitigate this, we exploit the inequality $\sqrt{x}\leq x$ for $x\geq 1$, allowing us to upper bound the regularizer as:
@@
WDPO Training: (1) Emotion alignment: The model was trained for 40 epochs with an effective batch size of 64. We used Adam optimizer, with a learning rate of $5.0\times 10^{-7}$ following 12 warmup steps. A maximum gradient norm of 10 was applied to ensure stable training. The DPO beta parameter was set to 0.1 for all training runs. Experiments were conducted on a single 40 GB A100 GPU, requiring gradient accumulation over two steps. (2) LLaMA experiments: The models were trained for 8 epochs with an effective batch size of 128. We used Adam optimizer with a learning rate of $5.0\times 10^{-7}$ after $10\%$ warmup ratio and then the learning rate was reduced using a cosine scheduler. The DPO beta parameter was set to 0.01 for all training runs. Experiments were conducted on an 8xH100 GPU setup, requiring loading the model in bfloat16 and training with DeepSpeed ZeRO-2 optimizer (Rajbhandari et al.,, 2020).
KLDPO Implementation: In line 3 of the tractable KLDPO algorithm (Algorithm 2), we compute the approximate worst-case kernel $\underline{\mathsf{P}}(i)\propto\mathrm{exp}\left((1/\tau)(l(z_{i};\theta)-(1/n)\textstyle\sum\nolimits_{i=1}^{n}l(z_{i};\theta))\right)$. A key implementation challenge arises in distributed LLM training. A naive approach would calculate the $(1/n)\textstyle\sum\nolimits_{i=1}^{n}l(z_{i};\theta)$ term by averaging $l(z_{i};\theta)$ across all samples in the micro-batch of its respective worker. However, because micro-batch sizes are typically small in large-scale LLM training, this results in averaging over only a few samples, making the worst-case kernel highly noisy. To mitigate this, we introduce a synchronization step that performs an all-gather operation to collect $l(z_{i};\theta)$ values from all workers. This enables averaging over the full batch across all workers, significantly reducing the noise in the worst-case kernel.
KLDPO Training: (1) Emotion alignment: The model was trained for 40 epochs with an effective batch size of 64. We used Adam optimizer (Kingma and Ba,, 2014), with a learning rate of $5.0\times 10^{-7}$ following 12 warmup steps. A maximum gradient norm of 10 was applied to ensure stable training. The DPO beta parameter was set to 0.1 for all training runs. Experiments were conducted on a single 40 GB A100 GPU and gradient was accumulated over two steps to keep training consistent across all algorithms. (2) LLaMA experiments: The models were trained for 8 epochs with an effective batch size of 128. We used Adam optimizer with a learning rate of $5.0\times 10^{-7}$ after $10\%$ warmup ratio and then the learning rate was reduced using a cosine scheduler. The DPO beta parameter was set to 0.01 for all training runs. Experiments were conducted on an 8xH100 GPU setup, requiring loading the model in bfloat16 and training with DeepSpeed ZeRO-2 optimizer (Rajbhandari et al.,, 2020).
@@ Appendix G Limitations
Theoretical Limitations: Our theoretical analysis relies on Assumption 2, which ensures sufficient data coverage to guarantee strong convexity conditions. Although such data-coverage assumptions are standard within offline learning or fixed-dataset scenarios, they are moderately restrictive, as they require the training dataset to sufficiently cover the space of feature differences between the preferred and dis-preferred actions. The log-linear policy class assumption, while standard and easily extendable to neural network policies under mild additional conditions, does not constitute a significant limitation.
Experimental Limitations: Empirically, Wasserstein Direct Preference Optimization (WDPO) involves two separate gradient computations during training—one for calculating the gradient penalty and another for updating policy parameters via standard gradient descent. This dual-gradient requirement can increase computational complexity and training difficulty, potentially limiting practical scalability and efficiency compared to methods with a single gradient computation.

### v3

@@ Appendix C Proof of KLDPO Sample Complexity
Note that the function $g(\theta)=\mathbb{E}_{z\sim\mathsf{P}}[l(z;\theta)]$ is $\gamma$-strongly convex with respect to $\lVert\cdot\rVert_{\Sigma_{\mathsf{P}}}$ by Lemma˜10. We use this fact in $(a)$. The inequality in $(b)$ is due to $\sup_{x}(f(x)+g(x))\leq\sup_{x}f(x)+\sup_{x}g(x)$. The last inequality $(c)$ is because $\operatorname{\lambda_{\mathrm{min}}}(\Sigma_{\mathsf{P}})\geq\lambda$, for all $\mathsf{P}\in\mathcal{P}_{\mathrm{KL}}$ by Assumption˜2. This implies that $\mathcal{L}^{\mathrm{KL}}$ is a $\gamma\lambda$-strongly convex function with respect to $\lVert\cdot\rVert_{2}$. ∎
@@ C.1 Proof of Policy Parameter Convergence of KLDPO
By Lemma˜13, we have that, with probability at least $1-\delta$,
@@
where the first inequality is because $\theta^{\mathrm{KL}}_{n}$ is the minimizer of $\mathcal{L}^{\mathrm{KL}}_{n}$. Now by the $\gamma\lambda$-strong convexity of $\mathcal{L}^{\mathrm{KL}}$ (see Lemma˜14) and Lemma˜5.II, we have that
@@ Appendix D Proof of Tractable KLDPO
Next, we prove the formal version of Proposition˜2.
@@
Here, $\mu$ and $\lambda$ are implicitly defined by the constraints (Eq.˜24-Eq.˜26). Now we prove that the dual variables $\mu-\lambda$ can be upper bounded by $\sum_{i=1}^{n}q_{i}l(z_{i};\theta)$.
@@
$\mu-\lambda$ satisfies $\mu-\lambda\leq-\sum_{i=1}^{n}\mathsf{P}_{n}^{o}(i)l(z_{i};\theta)$.
@@
This implies that $\mu-\lambda\leq-\sum_{i=1}^{n}q_{i}l(z_{i};\theta)$. ∎
@@ E.1 ArmoRM Multi-objective Alignment
Similar to the Emotion Alignment experiments, we generate preference labels according to convex combinations of two reward objectives, i.e., $r^{*}_{\mathrm{convex}}$ defined as Mixture Evaluation in previous section. Specifically, we consider three pairs of objectives: (1) Ultrafeedback-Honesty and Helpsteer-Complexity, (2) Ultrafeedback-Helpfulness and Helpsteer-Coherence, and (3) Ultrafeedback-truthfulness and Helpsteer-Complexity. We generate preference labels according to $\alpha^{o}=0.5$ for all three cases. All models are trained for 4 epochs. Then we introduce reward shift by evaluating WDPO, KLDPO, and DPO on $r^{*}_{\mathrm{convex}}(\alpha)$, where $\alpha\in\{0,0.1,0.3,0.5,0.7,0.9,1\}$. In the first plot of Fig.˜4, the training preferences are generated accoding to reward pair (1). We observe that WDPO and KLDPO achieve superior performance compared to DPO. In particular, when reward distribution shift happens in two directions (towards standalone Ultrafeedback-Honesty and Helpsteer-Complexity), they clearly outperform DPO. In the middle plot, the training preferences are generated according to reward pair (2). We observe that WDPO with both $\rho_{o}=0.005$ and $\rho_{o}=0.01$ are particularly robust against reward distribution shift. Lastly, in the third plot, the training preferences are generated according reward pair (3). We observe that both WDPO and KLDPO achieve notable robustness compared to DPO.
@@ E.2 Leaderboard Alignment
In this section, we include alignment results evaluated on all OpenLLM Leaderboard v2 (Fourrier et al., 2024) sub-tasks. We list all sub-task names in Table˜2.
Table 2: All sub-task names in OpenLLM Leaderboard v2.
@@
LLaMA-3.2-1B results: In Table˜3, we compare DPO, KLDPO, and WDPO trained using LLaMA-3.2-1B on all 39 sub-tasks of OpenLLM Leaderboard v2. We observe that our WDPO and KLDPO methods achieve superior alignment performance on the majority of subtasks. Although WDPO and KLDPO slightly underperform on few subtasks, their primary strength lies in generalization, precisely because they consistently enhance performance across a diverse range of subtasks.
LLaMA-3.1-8B results: In Table˜4, we compare DPO and KLDPO, both trained using LLaMA-3.1-8B. Earlier, we demonstrated that WDPO, trained on LLaMA-3.2-1B, outperforms both DPO and KLDPO. However, WDPO’s requirement for dual gradient computations increases its computational complexity. Due to resource constraints, we present KLDPO results for the 8B model, as it is more scalable. Following the LLaMA-3.2-1B experiments, we train KLDPO for two epochs, the point where DPO achieved optimal robustness via early stopping. Notably, KLDPO exhibits exceptional performance on math-related tasks.
@@
Table 3: Evaluation of DPO, KLDPO, and WDPO on all OpenLLM Leaderboard v2 sub-tasks.
@@
Table 4: Evaluation of DPO and KLDPO on all OpenLLM Leaderboard v2 sub-tasks.
@@ Appendix F Additional Experiment Details
Reward Model Training: The raw Emotion dataset (Saravia et al., 2018) consists of text samples paired with multi-class labels for six different emotion classes (joy, sadness, love, anger, fear, and surprise). This dataset was then transformed into a multi-label dataset, referred to as the Emotion Reward Dataset. To create the multi-label dataset, the surprise class was excluded due to its limited representation in the original dataset. Following this, up to three random text samples from the raw dataset were concatenated, and their associated labels were merged. This pre-processing step ensured that the reward model encountered text samples representing multiple emotions during training.
@@
Supervised Fine-Tuning: Before training the WDPO algorithm, it is essential to ensure that the model familiarize with the types of texts present in the dataset. To achieve this, we performed supervised fine-tuning (SFT). We selected GPT-2 as the base language model and trained it to predict the next token based on the text samples in the emotion dataset. The maximum length of each text sample was capped at 68 tokens. The model was trained for 10 epochs with a batch size of 64. The training used the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $5.0\times 10^{-7}$ following 12 warmup steps. Additionally, a maximum gradient norm of 10 was applied to stabilize the training. The ArmoRM setups did not need any SFT as we used Intruct models which have already undergone multiple rounds of SFT and alignment.
@@
WDPO Implementation: (1) In WDPO training, one of the main challenges is calculating the gradient penalty of the DPO loss with respect to the input. However, since the input is tokenized as integers, gradient cannot be directly calculated. To address this, gradient is calculated with respect to the output of the embedding layer, where gradients are tracked. (2) In line 4 of the tractable WDPO algorithm (Algorithm˜1), we compute the gradient regularizer: $\mathcal{R}(\pi_{\theta};\mathcal{D})=\rho_{o}(\mathbb{E}_{z\sim\mathcal{D}}\lVert\nabla_{z}l(z;\theta)\rVert_{2}^{2})^{1/2}$. A key implementation challenge arises in distributed LLM training. A naive approach computes the gradient of the pointwise DPO loss with respect to each input, averages the gradient norms over the micro-batch, and applies this as a regularizer to the batch DPO loss on each worker. However, due to the typically small micro-batch sizes in large-scale LLM training, this averaging is performed over very few samples, resulting in a highly noisy and unstable gradient penalty. To mitigate this, we exploit the inequality $\sqrt{x}\leq x$ for $x\geq 1$, allowing us to upper bound the regularizer as:
@@
WDPO Training: (1) Emotion alignment: The model was trained for 40 epochs with an effective batch size of 64. We used Adam optimizer, with a learning rate of $5.0\times 10^{-7}$ following 12 warmup steps. A maximum gradient norm of 10 was applied to ensure stable training. The DPO beta parameter was set to 0.1 for all training runs. Experiments were conducted on a single 40 GB A100 GPU, requiring gradient accumulation over two steps. (2) LLaMA experiments: The models were trained for 8 epochs with an effective batch size of 128. We used Adam optimizer with a learning rate of $5.0\times 10^{-7}$ after $10\%$ warmup ratio and then the learning rate was reduced using a cosine scheduler. The DPO beta parameter was set to 0.01 for all training runs. Experiments were conducted on an 8xH100 GPU setup, requiring loading the model in bfloat16 and training with DeepSpeed ZeRO-2 optimizer (Rajbhandari et al., 2020).
KLDPO Implementation: In line 3 of the tractable KLDPO algorithm (Algorithm˜2), we compute the approximate worst-case kernel $\underline{\mathsf{P}}(i)\propto\mathrm{exp}\left((1/\tau)(l(z_{i};\theta)-(1/n)\textstyle\sum\nolimits_{i=1}^{n}l(z_{i};\theta))\right)$. A key implementation challenge arises in distributed LLM training. A naive approach would calculate the $(1/n)\textstyle\sum\nolimits_{i=1}^{n}l(z_{i};\theta)$ term by averaging $l(z_{i};\theta)$ across all samples in the micro-batch of its respective worker. However, because micro-batch sizes are typically small in large-scale LLM training, this results in averaging over only a few samples, making the worst-case kernel highly noisy. To mitigate this, we introduce a synchronization step that performs an all-gather operation to collect $l(z_{i};\theta)$ values from all workers. This enables averaging over the full batch across all workers, significantly reducing the noise in the worst-case kernel.
KLDPO Training: (1) Emotion alignment: The model was trained for 40 epochs with an effective batch size of 64. We used Adam optimizer (Kingma and Ba, 2014), with a learning rate of $5.0\times 10^{-7}$ following 12 warmup steps. A maximum gradient norm of 10 was applied to ensure stable training. The DPO beta parameter was set to 0.1 for all training runs. Experiments were conducted on a single 40 GB A100 GPU and gradient was accumulated over two steps to keep training consistent across all algorithms. (2) LLaMA experiments: The models were trained for 8 epochs with an effective batch size of 128. We used Adam optimizer with a learning rate of $5.0\times 10^{-7}$ after $10\%$ warmup ratio and then the learning rate was reduced using a cosine scheduler. The DPO beta parameter was set to 0.01 for all training runs. Experiments were conducted on an 8xH100 GPU setup, requiring loading the model in bfloat16 and training with DeepSpeed ZeRO-2 optimizer (Rajbhandari et al., 2020).
@@ Appendix G Limitations
Theoretical Limitations: Our theoretical analysis relies on Assumption˜2, which ensures sufficient data coverage to guarantee strong convexity conditions. Although such data-coverage assumptions are standard within offline learning or fixed-dataset scenarios, they are moderately restrictive, as they require the training dataset to sufficiently cover the space of feature differences between the preferred and dis-preferred actions. The log-linear policy class assumption, while standard and easily extendable to neural network policies under mild additional conditions, does not constitute a significant limitation.
Experimental Limitations: Empirically, Wasserstein Direct Preference Optimization (WDPO) involves two separate gradient computations during training, one for calculating the gradient penalty and another for updating policy parameters via standard gradient descent. This dual-gradient requirement can increase computational complexity and training difficulty, potentially limiting practical scalability and efficiency compared to methods with a single gradient computation.